Calculate the number of protons neutrons , electrons and neutrons of calcium ion of CaCl2?
1
Expert's answer
2017-02-21T09:34:08-0500
The ionic charge tells the number of electrons gained or lost. Since electrons are negative, adding an electron will make the ion have a negative charge, and losing electrons will make for a positively charged ion.
In CaCl2 Ca is in ion form.
A calcium(II) ion, Ca2+, has 20 protons and 18 electrons because it has lost two (to give it the +2 charge).
p( Ca2+) = 20
e ( Ca2+) =20 - 2 = 18
n( Ca2+) = Ar ( Ca) - p( Ca2+)
n( Ca2+) = 40 - 20 = 20